find the particular solution and general solution of the equation y''''+y'''=e^(2x)

[25] Find a particular solution and the general solution of the equation y(4) + y = 220
Answer #1

(04+03)y = epX Where Doelen Auxiliary equation r28220 = 888+1)=0 r= 0,0,0,-) Y = c + Cq X + Cq x² + Cu ex Now find the partic